# Gaussian process regression with additive periodic kernels for two-body interaction analysis in coupled phase oscillators

### v1

@@ Abstract
We propose a Gaussian process regression framework with additive periodic kernels for the analysis of two-body interactions in coupled oscillator systems. While finite-order Fourier expansions determined by Bayesian methods can still yield artifacts such as a high-amplitude, high-frequency vibration, our additive periodic kernel approach has been demonstrated to effectively circumvent these issues. Furthermore, by exploiting the additive and periodic nature of the coupling functions, we significantly reduce the effective dimensionality of the inference problem. We first validate our method on simple coupled phase oscillators and demonstrate its robustness to more complex systems, including Van der Pol and FitzHugh-Nagumo oscillators, under conditions of biased or limited data. We next apply our approach to spiking neural networks modeled by Hodgkin-Huxley equations, in which we successfully recover the underlying interaction functions. These results highlight the flexibility and stability of Gaussian process regression in capturing nonlinear, periodic interactions in oscillator networks. Our framework provides a practical alternative to conventional methods, enabling data-driven studies of synchronized rhythmic systems across physics, biology, and engineering.
@@ 1 Introduction
The coupled phase-oscillator model [1, 2, 3] is commonly used to describe rhythmic behavior of various complex systems, such as the coordinated activity of neurons in the brain [4] or the synchronized flashing of fireflies [5]. This model allows for significant dimensionality reduction and can provide insight into the rhythmic behavior of the original system.
The coupling functions between the oscillators are represented as functions of the phase difference, which determine the synchronisation characteristics of the system. In other words, estimating the coupling functions from observed data is a crucial step in revealing the underlying mechanisms of rhythmic phenomena. Many studies have addressed to this issue using methods such as Fourier series expansion [6, 7, 8, 9]. However, the truncation order of Fourier series, which determines the complexity of the model and significantly affects the approximation accuracy, is typically arbitrary and subject to the limitations of the evidence approximation. Furthermore, the estimated Fourier coefficients may take large absolute values, leading to a high-amplitude, high-frequency oscillation. This phenomenon is undesirable since it means that coupling functions oscillate at unnaturally high frequencies.
To overcome these difficulties associated with Fourier truncation, we use the Gaussian process regression, which is a non-parametric regression method that allows for a flexible modeling of complex functions through the use of a covariance function [10]. Specifically, we propose a novel approach that incorporates key characteristics of the coupling functions—periodicity and additivity—into the kernel, introducing a periodic additive kernel.
@@
This paper is organized as follows. In section 2, we briefly review coupled phase-oscillator models and Gaussian process. In section 3, we propose a method for extracting the coupling functions of oscillator systems from data exhibiting rhythmic phenomena using Gaussian process regression. This allows for more flexible estimation than exsiting methods. The proposed method is applied to various systems in the next section 4. In particular, we demonstrate the superiority of the proposed method over existing methods using coupled FitzHugh-Nagumo oscillators, under conditions of biased or limited data. Furthermore, the proposed method is applied to a many-body system of spiking neurons and it is confirmed that all coupling functions are estimated with high accuracy. Finally, we conclude this paper and provide further remarks in Section 5.
@@ 2.1 Coupled phase oscillators
Here, the function $F_{i}$ represents the spontaneous dynamics of the element $i$, and $G_{j\to i}$ represents a weak action from the element $i$ to the element $j$. We assume that each element has a stable limit cycle (i.e., each element is an oscillator). It is well known that such a system can be reduced to coupled phase oscillators [1, 2].
@@
The functions $\{\Gamma_{ij}\}$ are coupling functions, which represent the action from the oscillator $j$ to the oscillator $i$.
The advantage of employing phase-oscillator model lies in the significant simplification of the system. While the original system (1) is generally multidimensional, the phase oscillators (2) assign only one variable $\phi_{i}$ to each limit cycle. Consequently, the focus can be directed solely towards rhythm synchronization characteristics.
2.2 estimation for coupling functinos
In this paper, we estimate all phase coupling functions $\{\Gamma_{ij}\}$ from phase time series $\mathcal{D}=\{(\bm{x}_{j},y_{j})\}_{j=1}^{n_{\mathrm{data}}}$ with $\bm{x}_{j}\in\mathbb{T}^{N-1}$ and $y_{j}\in\mathbb{R}$ for $j=1,2,\ldots,n_{\mathrm{data}}$. Here, $\mathbb{T}$ represents a torus. Since the phase coupling functions determine synchronization characteristics of the oscillators, it is desirable to estimate these functions from observed time series.
@@ 3 Methodology: Gaussian process regression
Suppose we have a time-series data $\mathcal{D}=\{(\bm{x}_{j},y_{j})\}_{j=1}^{n_{\mathrm{data}}}\subset\mathbb{T}^{N-1}\times\mathbb{R}$ with $n_{\mathrm{data}}$ number of data and a function $f\colon\mathbb{T}^{N-1}\to\mathbb{R}$, such that
$y_{i}=f(\bm{x}_{i})+\xi_{i},$ (3)
for $i=1,\dots,n_{\mathrm{data}}$, where $\xi_{i}$ is a random variable of $\mathcal{N}(0,\sigma^{2})$ that represents a noise in the output. Our task is to predict the unknown function $f$ from the data $\mathcal{D}$ using the Gaussian process.
We start by assuming that the unknown regression function $f$ is drawn from a given Gaussian process prior,
$f\sim\mathcal{GP}(m,k),$ (4)
where $m\colon\mathbb{T}^{N-1}\to\mathbb{R}$ is called a mean function and $k\colon\mathbb{T}^{N-1}\times\mathbb{T}^{N-1}\to\mathbb{R}$ is called a covariance function. The covariance function should be chosen so that it reflects one’s prior knowledge or belief about the regression function $f$, and we will discuss this in the next subsection. In many cases, the mean function $m$ is set to a constant zero function for simplicity.
The posterior distribution of $f$ is calculated analytically by linear algebra, and is again a Gaussian process. The distribution is given by the following closed form:
$f\mid\mathcal{D}\sim\mathcal{GP}(\overline{m},\overline{k}),$ (5)
where the posterior mean function $\overline{m}$ and the posterior covariance function $\overline{k}$ are
$\overline{m}(\bm{x})$ $=m(\bm{x})+k_{\bm{x}X}(k_{XX}+\sigma^{2}I_{n_{\mathrm{data}}})^{-1}(\bm{y}-m_{X}),$ (6)
$\overline{k}(\bm{x},\bm{x}^{\prime})$ $=k(\bm{x},\bm{x}^{\prime})-k_{\bm{x}X}(k_{XX}+\sigma^{2}I_{n_{\mathrm{data}}})^{-1}k_{X\bm{x}^{\prime}}.$ (7)
@@
$[k_{XX}]_{ij}$ $=k(\bm{x}_{i},\bm{x}_{j}),$
$k_{X\bm{x}}$ $=k_{\bm{x}X}^{\top}=(k(\bm{x}_{1},\bm{x}),\dots,k(\bm{x}_{n_{\mathrm{data}}},\bm{x}))^{\top},$
$m_{X}$ $=(m(\bm{x}_{1}),\dots,m(\bm{x}_{n_{\mathrm{data}}}))^{\top},$
$\bm{y}$ $=(y_{1},\dots,y_{n_{\mathrm{data}}})^{\top}.$
@@ 3.1 Choice of the covariance function
The covariance function $k$ is a fundamental component of Gaussian process regression, as it encodes our assumptions about the function which we wish to predict (see [11] for a detailed discussion). In our problem setting, since the coupling function have a domain of torus $\mathbb{T}^{N-1}$, which means the function is $2\pi$-periodic, it is appropriate to use a periodic kernel as the covariance function. In addition, since the coupling function is additive (i.e., it is represented as the sum of two-body interactions), it can be seen that the kernel is also additive [12]. In conclusion, we propose to use following additive periodic kernel to estimate the coupling functions:
$k(\bm{x},\bm{x}^{\prime})=\sum_{j=1}^{N-1}\theta_{0}^{(j)}\exp\left(\theta_{1}^{(j)}\cos(x_{j}-x_{j}^{\prime})\right).$ (8)
@@ 3.2 Optimization
$\begin{split}\mathcal{L}(\bm{\theta})&:=\log p(\bm{y}\mid X,\bm{\theta})\\ &=-\frac{1}{2}\bm{y}^{\top}K_{\bm{\theta}}\bm{y}-\frac{1}{2}\log\det K_{\bm{\theta}}-\frac{n_{\mathrm{data}}}{2}\log 2\pi,\end{split}$ (9)
where $K_{\bm{\theta}}=k_{XX}+\sigma^{2}I_{n_{\mathrm{data}}}$ is the covariance matrix for the noisy data $\mathcal{D}$. The most probable parameters $\bm{\theta}$ are, therefore, estimated by finding the maximum of the log-likelihood function $\mathcal{L}(\bm{\theta})$.
@@
Stochastic gradient decent (SGD) method, regarded as a stochastic approximation of gradient descent optimization methods, is another way of optimization. In the SGD method, the gradient is determined by different minibatches at each step, and parameter updating proceeds accordingly. The greatest benefit is that it requires less computation time to converge than the gradient method, which uses all data at each step. It is also thought to be less likely to be trapped in local solutions, thanks to stochastic fluctuations in the gradient direction for each minibatch. Note that we can calculate the derivative of the covariance function analytically with respect to their hyperparameters. This enables efficient optimisation using automatic differentiation. The implementation was carried out using GPflow [13].
3.3 The significance of using Gaussian process regression
One of the main merits of Gaussian process regression is its ability to provide uncertainty estimates for the regression parameters, which can be useful in situations where the data are noisy or sparse. It also allows for the modeling of complex functions without requiring the specification of a fixed functional form, which can be difficult to determine in many cases. Additionally, the use of a kernel function allows for the incorporation of prior knowledge about the function, such as its smoothness or periodicity, which can improve the accuracy of the estimates.
Gaussian process regression has the advantage that all calculations can be done in a closed form using only matrix operations. However, as the number of data $n_{\mathrm{data}}$ increases, inverse matrix calculations for $n_{\mathrm{data}}\times n_{\mathrm{data}}$ matrices are required, and the amount of memory and computation is enormous. Many sparse approximations have been proposed to overcome the computational complexity of Gaussian process regression [14, 15, 16, 17, 18].
@@ 4.1 Coupled phase oscillators
[…]
where $\phi_{1},\phi_{2},\phi_{3}$ are the phases of the oscillators. We prepare the time series by solving the ODEs (11) with forward Euler method from $t_{0}=0.0$ to $t_{1}=15.0$ using a randomly selected initial value and sample the data with sampling rate 0.2. We repeat the above calculation 50 times. Estimation results are shown in Figure 1.
@@ 4.2 Coupled Van der Pol oscillators
where $\hat{x}_{i}$ and $\hat{y}_{i}(t)$ are centered time series of $x_{i}(t)$ and $y_{i}(t)$, respectively. Even though $\vartheta_{i}(t)\in[0,2\pi)$ for all $t$, it is inappropriate to consider $\vartheta_{i}(t)$ as the phase time series of the oscillator $i$. This is because $\vartheta_{i}$ does not vary monotonically in the absence of interactions and noise. To overcome this difficulty, the following transformation from $\vartheta_{i}$ to $\theta_{i}$ should suffice [19, 20]:
@@
Estimation results are shown in Figure 2. The results are highly consistent with the true coupling functions derived using the adjoint method. For comparison purposes, we also present the estimated coupling functions with the existing method [9].

### v2

# arXiv:2504.09110v2

@@ Abstract
Since many physical laws—from classical mechanics to electromagnetism—are formulated as two-body interactions, the same perspective naturally extends to biological and social dynamics. Here we focus on rhythmic phenomena, where phase reduction theory shows that synchronization dynamics can be universally described by coupled phase oscillators. Estimating the interaction functions of such systems from data offers a direct path to understanding and predicting such dynamics. Existing Fourier-series-based methods encounter difficulties with limited or biased data. To overcome this, we employ Gaussian process regression with additive periodic kernels. In our approach, we incorporate information about the estimation target into the statistical model in advance by designing kernel functions that capture the characteristics—additivity and $2\pi$-periodicity—of the coupling functions. Furthermore, owing to the Bayesian framework, our method enables the evaluation of uncertainty in the estimation results. We validate our approach on Van der Pol, FitzHugh-Nagumo, and spiking neural models. Our approach outperforms Fourier-series baselines in both error and stability under biased phase sampling. This enables data-driven studies of rhythm dynamics across a broader range of datasets. Furthermore, it makes a first step toward a statistically grounded, data-driven approach to general many-body systems with two-body interactions.
[…]
A55
@@ 1 Introduction
Figure 1: Conceptual overview. (Left) Not only basic physical laws such as classical mechanics, but also interactions in biological networks and rhythm synchronization can be formulated as the sum of two-body interactions. Rhythmic synchronization phenomena are universally described by coupled phase oscillators with a phase difference $\Delta\phi_{ij}$. (Right) The two-body coupling function is estimated by Gaussian process regression using an additive periodic kernel. Reflecting the phase bias reconstructed from the rhythmic time series, the uncertainty of the estimation results is expressed as the variance of the posterior distribution.
Interactions in many-body systems are often expressed as superpositions of two-body interactions (Fig. 1, Left). Typical examples include the gravitational force between point masses in classical mechanics and the Coulomb force between point charges in electromagnetism. Even when higher-order interactions exist in the system, it is meaningful to first approximate the system by considering only two-body interactions. This leads us to believe that, when analyzing unknown many-body systems such as communication among numerous organisms or active matter, it is reasonable to expect that the interactions between elements can be expressed as the sum of two-body interactions. If such interactions can be estimated from observational data, this provides a powerful means of uncovering underlying physical laws.
Numerous methods have been proposed for estimating interactions from observational data Galan2005 ; Tokuda2007 ; Stankovski2012 ; ota2014 ; Brunton2016 ; Raissi2018 ; Chen2018 ; Lu2021 . With recent advances in machine learning, it has become feasible to estimate even highly nonlinear interactions when sufficient data are available. In particular, kernel methods are effective for modeling nonlinear relationships and, when combined with Bayesian inference, enable quantification of the uncertainty in the estimated functions. However, to apply such approaches to the estimation of two-body interaction functions, the kernel must be expressible as a sum of two-body kernels. Generic kernels typically do not satisfy this decomposition, which limits their applicability in this context.
In this paper, we focus on rhythmic interactions, which are known to be reduced to systems of coupled phase oscillators Kuramoto1984 ; Winfree1980 ; Nakao2016 . This model has recently gained renewed prominence, as highlighted by the 2025 Boltzmann Medal awarded to Professor Yoshiki Kuramoto for his foundational contributions to the statistical physics of synchronization. Owing to its mathematical simplicity and universality, this model is widely employed to describe the rhythmic behavior of diverse complex systems, such as the coordinated activity of neurons in the brain Viriyopase2018 or the synchronized flashing of fireflies Ermentrout1991 . However, existing methods for estimating rhythm interactions ota2014 based on Fourier series face difficulties in accurately estimating the interaction function of oscillators in a synchronized state, due to the bias in the data. To overcome this difficulty, we employ Gaussian process regression, a nonparametric kernel method Rasmussen2006 , and propose a new approach that embeds the essential features of coupling functions in phase-oscillator models—namely periodicity and additivity—into the kernel design, introducing an additive periodic kernel (Fig. 1, Right).
@@
This paper is organized as follows. In Section 2, we briefly review coupled phase-oscillator models. In Section 3, we propose a method for extracting the coupling functions of oscillator systems from data exhibiting rhythmic phenomena using Gaussian process regression. The proposed method is then applied to various systems in Section 4. In particular, we demonstrate the superiority of our approach over existing methods using coupled FitzHugh–Nagumo oscillators under conditions of biased or limited data. Furthermore, we apply the method to a many-body system of spiking neurons and confirm that all coupling functions are estimated with high accuracy. Finally, we conclude the paper and discuss further applications in Section 5.
@@ 2.1 Coupled phase oscillators
Here, the function $F_{i}$ represents the spontaneous dynamics of the element $i$, and $G_{j\to i}$ represents a weak action from the element $i$ to the element $j$. We assume that each element has a stable limit cycle (i.e., each element is an oscillator). It is well known that such a system can be reduced to coupled phase oscillators Kuramoto1984 ; Winfree1980 .
@@
The functions $\{\Gamma_{ij}\}$ are called coupling functions, and for each $i$ and $j$, $\Gamma_{ij}$ represents the action from the oscillator $j$ to the oscillator $i$.
The advantage of employing the phase-oscillator model lies in the significant simplification of the system. While the original system (1) is generally multidimensional, the phase oscillators (2) assign only one variable $\phi_{i}$ to each limit cycle. Consequently, the focus can be directed solely towards rhythm synchronization characteristics.
2.2 estimation for coupling functions
In this paper, we estimate all phase coupling functions $\{\Gamma_{ij}\}$ from sampled phase time series $\{\bm{\phi}(\ell\delta t)\}_{\ell=1}^{n_{\mathrm{data}}}$ with $\bm{\phi}(\ell\delta t):=(\phi_{1}(\ell\delta t),\cdots,\phi_{N}(\ell\delta t))\in\mathbb{T}^{N}$ for $\ell=1,2,\ldots,n_{\mathrm{data}}$. Here, $\delta t$ denotes the sampling interval and $\mathbb{T}:=[0,2\pi)$. Since the phase coupling functions determine synchronization characteristics of the oscillators, it is desirable to estimate these functions from observed time series.
@@ 3 Methodology: Gaussian process regression
Suppose we have a time-series data $\mathcal{D}=\{(\bm{\phi}(\ell\delta t),\bm{y}(\ell\delta t))\}_{\ell=1}^{n_{\mathrm{data}}}$, which
$\bm{y}(\ell\delta t)$ $:=\left(\frac{\phi_{1}(\ell\delta t+\delta t)-\phi_{1}(\ell\delta t)}{\delta t},\cdots,\frac{\phi_{N}(\ell\delta t+\delta t)-\phi_{N}(\ell\delta t)}{\delta t}\right)\in\mathbb{R}^{N},$ (3)
and a function $\Gamma_{ij}\colon[0,2\pi)\to\mathbb{R}$, such that
$\begin{split}y_{i}&=f_{i}(\Delta\bm{\phi}_{i})+\xi_{i},\quad f_{i}(\Delta\bm{\phi}_{i})=\omega_{i}+\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{N}\Gamma_{ij}(\Delta\phi_{ij}),\\ \Delta\bm{\phi}_{i}&:=(\Delta\phi_{i1},\cdots,\Delta\phi_{iN})\in\mathbb{T}^{N-1}\end{split}$ (4)
for $\ell=1,\dots,n_{\mathrm{data}}$, where $\xi_{i}$ is a random variable of $\mathcal{N}(0,\sigma^{2})$ that represents a noise in the output. Our task is to estimate the unknown functions $\{\Gamma_{ij}\}$ from the data $\mathcal{D}$ using the Gaussian process. If we estimate the sums $\{f_{i}\}$ of the two-body interactions, we can restore $\{\Gamma_{ij}\}$ except for the uncertainty of constants, so we will explain the estimation of $\{f_{i}\}$ below.
We start by assuming that the unknown regression function $f_{i}$ is drawn from a given Gaussian process prior,
$f_{i}\sim\mathcal{GP}(m_{i},k_{i}),$ (5)
where $m_{i}\colon\mathbb{T}^{N-1}\to\mathbb{R}$ is called a mean function and $k_{i}\colon\mathbb{T}^{N-1}\times\mathbb{T}^{N-1}\to\mathbb{R}$ is called a covariance function. The covariance function should be chosen so that it reflects one’s prior knowledge or belief about the regression function $f_{i}$, and we will discuss this in the next subsection. In many cases, the mean function $m_{i}$ is set to a constant zero function for simplicity.
The posterior distribution of $f_{i}$ is calculated analytically by linear algebra, and is again a Gaussian process. The distribution is given by the following closed form:
$f_{i}\mid\mathcal{D}\sim\mathcal{GP}(\overline{m}_{i},\overline{k}_{i}),$ (6)
where the posterior mean function $\overline{m}_{i}$ and the posterior covariance function $\overline{k}_{i}$ are
$\overline{m}_{i}(\Delta\bm{\psi})$ $=m_{i}(\bm{\psi})+k_{i;\bm{\psi}\Phi}(k_{i;\Phi\Phi}+\sigma^{2}I_{n_{\mathrm{data}}})^{-1}(\bm{y}_{i}-m_{i;\Phi}),$ (7)
$\overline{k}_{i}(\Delta\bm{\psi},\Delta\bm{\psi}^{\prime})$ $=k_{i}(\Delta\bm{\psi},\Delta\bm{\psi}^{\prime})-k_{i;\bm{\psi}\Phi}(k_{i;\Phi\Phi}+\sigma^{2}I_{n_{\mathrm{data}}})^{-1}k_{i;\Phi\bm{\psi}^{\prime}}.$ (8)
@@
$[k_{i;\Phi\Phi}]_{mn}$ $=k_{i}(\Delta\bm{\phi}_{i}(m\delta t),\Delta\bm{\phi}_{i}(n\delta t)),$
$k_{i;\Phi\bm{\psi}}$ $=k_{i;\bm{\psi}\Phi}^{\top}=(k_{i}(\Delta\bm{\phi}_{i}(0),\Delta\bm{\psi}),\dots,k(\Delta\bm{\phi}_{i}(n_{\mathrm{data}}\delta t),\Delta\bm{\psi}))^{\top},$
$m_{i;\Phi}$ $=(m_{i}(\Delta\bm{\phi}_{i}(0)),\dots,m_{i}(\Delta\bm{\phi}_{i}(n_{\mathrm{data}}\delta t)))^{\top},$
$\bm{y}_{i}$ $=(y(\Delta\bm{\phi}_{i}(0)),\dots,y(n_{\mathrm{data}}\delta t))^{\top}.$
Gaussian process regression has the advantage that all calculations can be done in a closed form using only matrix operations. However, as the number of data $n_{\mathrm{data}}$ increases, inverse matrix calculations for $n_{\mathrm{data}}\times n_{\mathrm{data}}$ matrices are required, and the amount of memory and computation is enormous. Many sparse approximations have been proposed to overcome the computational complexity of Gaussian process regression fitc ; Quinonero2005 ; titsias09 ; hensman13 ; hensman15 .
@@ 3.1 Choice of the covariance function
The covariance function is a fundamental component of Gaussian process regression, as it encodes our assumptions about the function which we wish to predict (see kanagawa2018 for a detailed discussion). In our problem setting, since the coupling function have a domain $\mathbb{T}^{N-1}$, which means the function is $2\pi$-periodic, it is appropriate to use a periodic kernel as the covariance function. In addition, since the coupling function is additive (i.e., it is represented as the sum of two-body interactions), it can be seen that the kernel is also additive Duvenaud2011 . In conclusion, we propose to use following additive periodic kernel to estimate the coupling functions:
$k_{i}(\Delta\bm{\psi},\Delta\bm{\psi}^{\prime})=\sum_{j=1}^{N-1}\theta_{0}^{(j)}\exp\left(\theta_{1}^{(j)}\cos(\Delta\psi_{j}-\Delta\psi_{j}^{\prime})\right).$ (9)
@@ 3.2 Optimization
$\begin{split}\mathcal{L}(\bm{\theta})&:=\log p(\bm{y}\mid\bm{\phi},\bm{\theta})\\ &=-\frac{1}{2}\bm{y}^{\top}K_{\bm{\theta}}\bm{y}-\frac{1}{2}\log\det K_{\bm{\theta}}-\frac{n_{\mathrm{data}}}{2}\log 2\pi,\end{split}$ (10)
where $K_{\bm{\theta}}=k_{\Phi\Phi}+\sigma^{2}I_{n_{\mathrm{data}}}$ is the covariance matrix for the noisy data $\mathcal{D}$. The most probable parameters $\bm{\theta}$ are, therefore, estimated by finding the maximum of the log-likelihood function $\mathcal{L}(\bm{\theta})$.
@@
The stochastic gradient descent (SGD) method, viewed as a stochastic approximation to standard gradient-based optimization, offers an alternative approach to optimization. In SGD, the gradient is estimated using a different minibatch of data at each iteration, and the parameters are updated accordingly. Its main advantage lies in reduced computational cost compared to full-batch gradient methods, which require evaluating the gradient using all data at every step. Moreover, the inherent stochasticity in minibatch selection helps the algorithm avoid getting trapped in local minima. Furthermore, the derivatives of the covariance function with respect to its hyperparameters can be computed analytically, enabling efficient optimization via automatic differentiation. Note that the implementation was performed using GPflow GPflow2017 .
@@ 4.1 Coupled phase oscillators
$\begin{aligned} &\dot{\phi}_{1}=1.0+0.01\sin(\phi_{2}-\phi_{1})-0.02\sin(\phi_{3}-\phi_{1}),\\ &\dot{\phi}_{2}=1.1+0.04\sin(\phi_{1}-\phi_{2})+0.01\sin(\phi_{3}-\phi_{2}),\\ &\dot{\phi}_{3}=0.95+0.01\sin(\phi_{1}-\phi_{3})-0.03\sin(\phi_{2}-\phi_{3}),\end{aligned}$ (12)
where $\phi_{1},\phi_{2},\phi_{3}$ are the phases of the oscillators. We prepare the time series by solving the ODEs (12) with forward Euler method from $t_{0}=0.0$ to $t_{1}=15.0$ using a randomly selected initial value and sample the data with sampling rate 0.2. We repeat the above calculation 50 times. Estimation results are shown in Fig. 2.
@@ 4.2 Coupled Van der Pol oscillators
where $\hat{x}_{i}$ and $\hat{y}_{i}(t)$ are centered time series of $x_{i}(t)$ and $y_{i}(t)$, respectively. Even though $\vartheta_{i}(t)\in[0,2\pi)$ for all $t$, it is inappropriate to consider $\vartheta_{i}(t)$ as the phase time series of the oscillator $i$. This is because $\vartheta_{i}$ does not vary monotonically in the absence of interactions and noise. To overcome this difficulty, the following transformation from $\vartheta_{i}$ to $\theta_{i}$ should suffice Kralemann2007 ; Kralemann2008 :
@@
Estimation results are shown in Fig. 3. The results are highly consistent with the true coupling functions derived using the adjoint method. For comparison purposes, we also present the estimated coupling functions with the existing method ota2014 .